sammirosesharp:

Express answer in exact form. Show all work for full credit. A regular hexagon with sides of 3" is inscribed in a circle. Find the area of a segment formed by a side of the hexagon and the circle. (Hint: remember Corollary 1--the area of an equilateral triangle is 1/4s^2sqrt3)

Vocaloid:



Vocaloid:

first find area of circle - area of the hexagon to get the area of each of the outer segments, then divide that by 6 to find one segment as a hint the hexagon is made of 6 equilateral triangles with side 3 you are given the area of an equilateral triangle in terms of s

sammirosesharp:

So The area of the hexagon is 60 degrees... But I only see the 3 on the circle.. So would it be 60-3? I'm confused. I'm sorry. I am absolutely terrible at math, so this is really easily confusing for me.

Vocaloid:

the hexagon is made of 6 equilateral triangles the area of one equilateral triangle is given as 1/4s^2sqrt3 where s = 3 area of the entire hexagon = 6 * 1/4s^2sqrt3 where s = 3